How many premutations of the letters of the word " BAHRAIN" end with "R"?


1
Expert's answer
2021-03-31T13:56:23-0400

BAHRAIN

If R is the ending word,

Then, number of B= 1

number of A= 2

number of H= 1

number of I = 1

number of N= 1 left in the word "BAHRAIN"

Total number of letters = 1+2+1+1+1=6

So, Total number of permutations= "\dfrac{6!}{1!\times 2!\times1!\times1!\times1!}=\dfrac{6!}{2!}=360"